for the reactioN
NO4​   2NO2 ,at 350 K,the value of K=0.4.The value of Kp for the reaction at the same temperature would be
a)11.48 atm
b)1.148 atm
c)1.4 x10-2 atm
d)1.4x10-3 atm
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Use formula
 
Kp = Kc * (RT)^n
 
Where n = 2 -1 = 1
 
Hence
Kp = Kc * RT
 
 
Regards
Arun (askIITians forum expert)